OpenStudy (anonymous):

Help? Medal for best answer! Integral of (10+9sinx)/cosx

OpenStudy (anonymous):

redo is please it doesent look rite

OpenStudy (anonymous):

How does it not look right?

OpenStudy (anonymous):

wat is cosx

OpenStudy (anonymous):

What do you mean? It's the cosine of x, written as cosx

OpenStudy (anonymous):

The equation is written correctly...

OpenStudy (anonymous):

\[\frac{ 10+9x }{ x }\] like this

OpenStudy (anonymous):

No, it's like\[\int\limits_{}^{}\frac{ (10+9sinx) }{ cosx }\]